# boxedasideAside

### v1

@@ 3.2 Computing the dimensions of the $p$-rank strata of ¯$\overline{\mathcal{B}}_{E,g}$
Otherwise, the generic point of $W$ is represented by a smooth curve, meaning that $W=\overline{\Gamma}\subseteq\overline{\mathcal{M}}_{g}$ where $\Gamma$ is a component of $V_{f}(\mathcal{B}_{E,g})\subseteq\mathcal{M}_{g}$. Consider the Torelli morphism $j:\overline{\mathcal{M}}_{g}\to\tilde{\mathcal{A}}_{g}$ with $\tilde{\mathcal{A}}_{g}$ the compactification of the moduli space of principally polarized abelian varieties $\mathcal{A}_{g}$ as in [Ale04] and note that $\dim W=\dim j(W)$ by our assumption on $W$. Assume that $\dim j(W)\geq\dim j(W\cap\Delta)+2$. In a projective embedding of $\tilde{\mathcal{A}}_{g}$, we could then intersect $j(W)$ with $(\dim j(W)-1)$-hyperplanes in general position to get a complete $1$-dimensional subvariety $X$ that does not meet $j(W\cap\Delta)$. This $X$ would be a subvariety of $j(W)-j(W\cap\Delta)=j(\Gamma)$, whose preimage in ¯$\overline{\mathcal{M}}_{g}$ would thus be a complete $1$-dimensional subvariety of $\Gamma$. This is impossible since $\mathcal{B}_{E,g}$ and $\Gamma$ cannot contain such subvarieties by Remark 3.1. Therefore,
$\dim W\leq\dim j(W\cap\Delta)+1.$
Note that $\dim j(V_{f}(\delta_{1,g-2}))=g-3+f$ since the Torelli morphism ”forgets” the pair $q_{1},q_{2}\in H$ of points in the same orbit of hyperelliptic involution as in (1) of the proof of Lemma 3.10. Combining this with Lemmas 3.10 and 3.11, we conclude that $\dim j(W\cap\Delta)\leq g-3+f$, which implies $\dim W\leq g-2+f$. This finishes the proof. ∎
@@
    If $E$ is a supersingular elliptic curve, then for any $10\leq f\leq g-$ 1, there exists a smooth curve $D$ of genus $g_{D}=g$ and p-rank $f_{D}=f$, which is a double cover of $E$, except in the case when $p=3$, $g=2$, and $f=0$.
@@
    Let $E$ be a supersingular elliptic curve. Then, for any $20\leq f\leq g-$ 2, there is a component of $V_{f}(\overline{\mathcal{B}}_{E,g})$ that consists entirely of singular curves.
@@
The third part follows from the proof of Theorem 3.12. Moreover, we can explicitly describe these components. Given a supersingular elliptic curve $(E,O_{E})$ in characteristic $p>0$, a component of $V_{f}(\overline{\mathcal{H}}_{g-2})$, for $g\geq 4$, corresponding to a family of stable hyperelliptic curves $H$, and two distinct points $q,q^{\prime}\in H$ in the same orbit under the hyperelliptic involution, a component of $V_{f}(\overline{\mathcal{B}}_{g})$ consists of isomorphism classes of stable curves of the form $(E,O_{E})\cup_{q=O_{E}}(H,q,q^{\prime})\cup_{q^{\prime}=O_{E}}(E,O_{E})$; note that the dimension of this family equals $\dim V_{f}(\overline{\mathcal{H}}_{g-2,1})=g-2+f=g-2+f-f_{E}$, as desired. For $g=2$, this follows from the proof of Lemma 3.6, while for $g=3$, it follows similarly as in the proof of Lemma 3.10 using the identification ¯$\overline{\mathcal{H}}_{1}\cong\overline{\mathcal{M}}_{1,1}$.
To establish the existence of smooth bielliptic curves with a prescribed $p$-rank, we employ a technique similar to the one from the proof of [Pri19, Theorem 6.4]. Let $p>3$ be a prime number and $E$ be an elliptic curve in characteristic $p$ whose $p$-rank equals $f_{E}$. By Lemma 3.6, for any $f_{E}\leq f^{\prime}\leq 1+f_{E}$, there is a smooth bielliptic curve $D$ of genus $2$ with $f_{D}=f^{\prime}$. By Lemma 3.5 and by comparing dimensions, there is a smooth hyperelliptic curve $H\to\mathbb{P}^{1}$ of genus $g-2$ whose $p$-rank equals $f-f^{\prime}$. Given ramification points $r\in D$ and $r^{\prime}\in H$ of the structure bielliptic and hyperelliptic morphisms of $D$ and $H$, respectively, consider the curve $(D,r)\cup_{r=r^{\prime}}(H,r^{\prime})$. Any irreducible family of such curves has dimension $(2-2+f^{\prime}-f_{E})+(g-3+f-f^{\prime})=g-3+f-f_{E}$, and is therefore contained in a component of $V_{f}(\overline{\mathcal{B}}_{E,g})$ that cannot consist entirely of isomorphism classes of singular curves, due to the description provided earlier and Theorem 3.12. Consequently, a generic member of this component is represented by a smooth bielliptic curve of genus $g$ with $p$-rank $f$, which is a double cover of $E$. A similar argument applies for $p=3$ and $f\neq 0$. In the remaining case, $p=3$, $f_{E}=0$, and $f=0$, by Remark 3.9 there are no smooth curves of genus $2$ that are double covers of $E$. However, by Lemma 4.2 below, there is a smooth curve $D^{\prime}$ of genus $3$ that is a double cover of $E$ and whose $3$-rank equals 00. To show the existence of a smooth curve of genus $g\geq 4$ with $3$-rank 00 that is a double cover of $E$, we now use the existence of $D^{\prime}$ and a smooth (hyper)elliptic curve of genus $g-3$ with $3$-rank 00 and a similar argument to the one we just presented for $p>3$. ∎
@@ 3.3 The $p$-rank stratification of the locus of bielliptic curves $\mathcal{B}_{g}$
Recall that a smooth curve $D$ of genus $g\geq 2$ is bielliptic if there exists an elliptic curve $E$ and a double cover $D\to E$. We denote the locus of bielliptic curves in $\mathcal{M}_{g}$ by $\mathcal{B}_{g}$ and note that $\dim\mathcal{B}_{g}=2g-2$. Given an elliptic curve $E$, $\mathcal{B}_{E,g}$ can be viewed as a subset of $\mathcal{B}_{g}$. To describe the closure of $\mathcal{B}_{g}$ in ¯$\overline{\mathcal{M}}_{g}$, we proceed similarly to the construction in Section 2. Namely, one first considers the locus ¯$\overline{\mathcal{B}}_{g}^{adm}$ of admissible stable bielliptic curves. This is a proper Deligne-Mumford stack in characteristic $p\neq 2$ equipped with morphisms ¯$\overline{\mathcal{B}}_{g}^{adm}\to\overline{\mathcal{M}}_{g,2g-2}$ and ¯$\overline{\mathcal{B}}_{g}^{adm}\to\overline{\mathcal{M}}_{1,2g-2}$ which, respectively, extract the source curve and the target curve. By [SvZ20, Theorem 3.7], they induce well-defined proper morphisms
$\Psi:\overline{\mathcal{B}}_{g}^{adm}\to\overline{\mathcal{M}}_{g}\quad\text{ and }\quad\Phi:\overline{\mathcal{B}}_{g}^{adm}\to\overline{\mathcal{M}}_{1,1}$
@@
obtained by forgetting the marked points and stabilizing. The locus of the admissible smooth bielliptic curves $\mathcal{B}_{g}^{adm}$ is dense in ¯$\overline{\mathcal{B}}_{g}^{adm}$ and $\Psi:\mathcal{B}_{g}^{adm}\to\mathcal{B}_{g}$ is a quasi-finite morphism obtained only by forgetting the admissible structure. It follows that ¯$\overline{\mathcal{B}}_{g}=\Psi(\overline{\mathcal{B}}_{g}^{adm})\subseteq\overline{\mathcal{M}}_{g}$, the locus of stable bielliptic curves of genus $g$, is the closure of $\mathcal{B}_{g}$ in ¯$\overline{\mathcal{M}}_{g}$.
Corollary 3.15 ($p>2$).
@@
For every $g\geq 2$ and $0\leq f\leq g$, the $p$-rank $\leq f$ locus of stable bielliptic curves $V_{f}(\overline{\mathcal{B}}_{g})$ is pure of dimension $g-2+f$.
@@
Let $W$ be a component of $V_{f}(\overline{\mathcal{B}}_{g})$. As before, we find that $\dim W\geq g-2+f$, using the purity of the $p$-rank stratification of ¯$\overline{\mathcal{M}}_{g}$. Consider the morphism $\Phi:\overline{\mathcal{B}}_{g}^{adm}\to{\overline{\mathcal{M}}}_{1,2g-2}\to{\overline{\mathcal{M}}}_{1,1}$ induced by $(D\to(E,P_{1},\ldots,P_{2g-2}))\mapsto(\overline{E},P_{1})$, with $(\overline{E},P_{1})$ the stabilization of $(E,P_{1})$. We break into the following two cases:
@@
    If the generic point of $W$ is represented by a smooth curve, let $W^{\prime}$ denote the preimage of $W$ under the map $\Psi:\overline{\mathcal{B}}_{g}^{adm}\to\overline{\mathcal{M}}_{g}$; note that $\dim W=\dim W^{\prime}$. There are two possibilities:
    1. 1.
      If $\Phi(W^{\prime})$ is a single point in $\mathcal{M}_{1,1}\subseteq\overline{\mathcal{M}}_{1,1}$ that corresponds to some (smooth) elliptic curve $E$, we claim that $E$ is supersingular. If $E$ were an ordinary curve, it would follow from Theorem 3.12 that $\dim W=g-3+f$. However, this leads to a contradiction since $W$ must satisfy $g-2+f\leq\dim W$. It follows that $W$ coincides with a component of $V_{f}(\overline{\mathcal{B}}_{E,g})$ with $f_{E}=0$, giving $\dim W=g-2+f$.
@@
      Otherwise, $\Phi(W^{\prime})=\overline{\mathcal{M}}_{1,1}$ and there must be an element in $W^{\prime}$ that maps to the isomorphism class of the singular elliptic curve $(\mathbb{P}^{1},\infty,0,1)/\{0=1\}$ under $\Phi$. The image of this element under $\Psi$ is in $W\cap\Delta_{0}$, which implies $W\cap\Delta_{0}\neq\o$. Since ¯$\overline{\mathcal{M}}_{g}$ is a smooth stack, to conclude $\dim W\leq g-2+f$ suffices to show that $\dim(W\cap\Delta_{0})\leq g-3+f$. By [SvZ20, Section 3.4] and similarly to our description from Section 2.1 and (1), the boundary of $\mathcal{B}_{g}$ in ¯$\overline{\mathcal{B}}_{g}$ is the closure of the loci consisting of double covers of $(E,O_{E})\cup_{O_{E}=\infty}(\mathbb{P}^{1},\infty)$ ramified at some points, where $E$ is a (possibly singular) curve of genus $1$. If we denote by $\Gamma_{s}$ the locus of the described double covers where $E$ is singular, similar computations as earlier in this section show that $\dim\Gamma_{s}\leq g-3+f$. On the other hand, using induction on $g$ as in the proof of Theorem 3.12 with the base case $g=2$ essentially covered by Lemma 3.6, we find that the locus $\Gamma_{ns}$ of the remaining curves in $W\cap\Delta_{0}$ has dimension $\dim\Gamma_{ns}\leq g-3+f$; compare this with (2). Therefore, $\dim(W\cap\Delta_{0})\leq g-3+f$, and thus $\dim W\leq g-2+f$.
  • $\cdot$
@@
    If $W$ is not generically represented by a smooth curve, then $W$ consists entirely of isomorphism classes of singular curves. By the discussion above and the inequality $\dim W\geq g-2+f$, we conclude that $W\not\subseteq\Delta_{0}$. A similar discussion shows that $W$ could only possibly be the closure of the locus consisting of isomorphism classes of double covers of curves of the form $(E,O_{E})\cup_{O_{E}=\infty}(\mathbb{P}^{1},\infty)$ ramified above $22g-$ 2 distinct points $P_{i}\in\mathbb{P}^{1}-\{\infty\}$, $21\leq i\leq 2g-$ 2, and unramified otherwise, where $E$ is a genus $1$ curve. In addition to this, it has to be $f_{E}=0$. In this case, $\dim W=g-2+f$ as we concluded in the proof of Theorem 3.12 and Corollary 3.14. In other cases, if would follow $\dim W<g-2+f$, which is impossible by our earlier assumption.

### v2

@@ 3.2 Computing the dimensions of the $p$-rank strata of $\overline{\mathcal{B}}_{E,g}$
In the remaining case, the generic point of $W$ is represented by a smooth curve. That means that there is a component $\Gamma$ of $V_{f}(\mathcal{B}_{E,g})\subseteq\mathcal{M}_{g}$ such that $W=\overline{\Gamma}\subseteq\overline{\mathcal{M}}_{g}$. In particular, note that $\dim\Gamma=\dim W\geq g-2+f\geq 1$, and that $W\cap\Delta\neq\varnothing$ because of Lemma 3.2. Moreover, since $\Gamma\subseteq\mathcal{M}_{g}$, it follows that $\dim\Gamma=\dim j(\Gamma)$, which consequently implies that $\dim W=\dim j(W)$, while $W\cap\Delta\neq\varnothing$ implies that $\dim j(W\cap\Delta)\geq 0$. Let us prove the following inequality
$\dim W\leq\dim j(W\cap\Delta)+1,$ (4)
which will give us the desired upper bound on $W$; we do this using a similar argument as in [Oor74, Section 1]. Assume that $\dim j(W)\geq\dim j(W\cap\Delta)+2$. In a projective embedding of $\tilde{\mathcal{A}}_{g}$, we could then intersect $j(W)$ with $(\dim j(W)-1)$-hyperplanes in general position to get a complete $1$-dimensional subvariety $X$ that does not meet $j(W\cap\Delta)$. This $X$ would be a subvariety of $j(W)-j(W\cap\Delta)=j(\Gamma)$, whose preimage in $\overline{\mathcal{M}}_{g}$ would thus be a complete $1$-dimensional subvariety of $\Gamma$. This is impossible since $\mathcal{B}_{E,g}$ and $\Gamma$ cannot contain such subvarieties by Lemma 3.2. Therefore, the inequality (4) holds. Note that $\dim j(V_{f}(\delta_{1,g-2}))=g-3+f$ since the Torelli morphism ”forgets” the pair $q_{1},q_{2}\in H$ of points in the same orbit of hyperelliptic involution as in (1) of the proof of Lemma 3.10. Combining this with Lemmas 3.10 and 3.11, we conclude that $\dim j(W\cap\Delta)\leq g-3+f$, which implies $\dim W\leq g-2+f$. This finishes the proof. ∎
@@
    If $E$ is a supersingular elliptic curve, then for any $0\leq f\leq g-1$, there exists a smooth curve $D$ of genus $g_{D}=g$ and p-rank $f_{D}=f$, which is a double cover of $E$, except in the case when $p=3$, $g=2$, and $f=0$.
@@
    Let $E$ be a supersingular elliptic curve. Then, for any $0\leq f\leq g-2$, there is a component of $V_{f}(\overline{\mathcal{B}}_{E,g})$ that consists entirely of singular curves.
@@
The third part follows from the proof of Theorem 3.12. Moreover, we can explicitly describe these components. Given a supersingular elliptic curve $(E,O_{E})$ in characteristic $p>0$, a component of $V_{f}(\overline{\mathcal{H}}_{g-2})$, for $g\geq 4$, corresponding to a family of stable hyperelliptic curves $H$, and two distinct points $q,q^{\prime}\in H$ in the same orbit under the hyperelliptic involution, a component of $V_{f}(\overline{\mathcal{B}}_{E,g})$ consists of isomorphism classes of stable curves of the form $(E,O_{E})\cup_{q=O_{E}}(H,q,q^{\prime})\cup_{q^{\prime}=O_{E}}(E,O_{E})$; note that the dimension of this family equals $\dim V_{f}(\overline{\mathcal{H}}_{g-2,1})=g-2+f=g-2+f-f_{E}$, as desired. For $g=2$, this follows from the proof of Lemma 3.6, while for $g=3$, it follows similarly as in the proof of Lemma 3.10 using the identification $\overline{\mathcal{H}}_{1}\cong\overline{\mathcal{M}}_{1,1}$.
To establish the existence of smooth bielliptic curves with a prescribed $p$-rank, we employ a technique similar to the one from the proof of [Pri19, Theorem 6.4]. Let $p>3$ be a prime number and $E$ be an elliptic curve in characteristic $p$ whose $p$-rank equals $f_{E}$. By Lemma 3.6, for any $f_{E}\leq f^{\prime}\leq 1+f_{E}$, there is a smooth bielliptic curve $D$ of genus $2$ with $f_{D}=f^{\prime}$. By Lemma 3.5 and by comparing dimensions, there is a smooth hyperelliptic curve $H\to\mathbb{P}^{1}$ of genus g2g-2 whose $p$-rank equals $f-f^{\prime}$. Given ramification points $r\in D$ and $r^{\prime}\in H$ of the structure bielliptic and hyperelliptic morphisms of $D$ and $H$, respectively, consider the curve $(D,r)\cup_{r=r^{\prime}}(H,r^{\prime})$. Any irreducible family of such curves has dimension $(2-2+f^{\prime}-f_{E})+(g-3+f-f^{\prime})=g-3+f-f_{E}$, and is therefore contained in a component of $V_{f}(\overline{\mathcal{B}}_{E,g})$ that cannot consist entirely of isomorphism classes of singular curves, due to the description provided earlier and Theorem 3.12. Consequently, a generic member of this component is represented by a smooth bielliptic curve of genus $g$ with $p$-rank $f$, which is a double cover of $E$. A similar argument applies for $p=3$ and $f\neq 0$. In the remaining case, $p=3$, $f_{E}=0$, and $f=0$, by Remark 3.9 there are no smooth curves of genus $2$ that are double covers of $E$. However, by Lemma 5.1 below, there is a smooth curve $D^{\prime}$ of genus $3$ that is a double cover of $E$ and whose $3$-rank equals 0. To show the existence of a smooth curve of genus $g\geq 4$ with $3$-rank 0 that is a double cover of $E$, we now use the existence of $D^{\prime}$ and a smooth (hyper)elliptic curve of genus g3g-3 with $3$-rank 0 and a similar argument to the one we just presented for $p>3$. ∎
@@ 3.3 The $p$-rank stratification of the locus of bielliptic curves $\mathcal{B}_{g}$
Recall that a smooth curve $D$ of genus $g\geq 2$ is bielliptic if there exists an elliptic curve $E$ and a double cover $D\to E$. The locus of bielliptic curves, $\mathcal{B}_{g}$, satisfies $\mathcal{B}_{E,g}\subset\mathcal{B}_{g}$ for any elliptic curve $E$, and $\dim\mathcal{B}_{g}=2g-2$. To describe the closure of $\mathcal{B}_{g}$ in $\overline{\mathcal{M}}_{g}$, we proceed similarly to the construction in Section 2. Namely, one first considers the locus $\overline{\mathcal{B}}_{g}^{\rm adm}$ of admissible stable bielliptic curves. This is a proper Deligne-Mumford stack in characteristic $p\neq 2$ equipped with morphisms $\overline{\mathcal{B}}_{g}^{\rm adm}\to\overline{\mathcal{M}}_{g,2g-2}$ and $\overline{\mathcal{B}}_{g}^{\rm adm}\to\overline{\mathcal{M}}_{1,2g-2}$ which, respectively, extract the source curve and the target curve. By [SvZ20, Theorem 3.7], they induce well-defined proper morphisms
$\Psi:\overline{\mathcal{B}}_{g}^{\rm adm}\to\overline{\mathcal{M}}_{g}\quad\text{ and }\quad\Phi:\overline{\mathcal{B}}_{g}^{\rm adm}\to\overline{\mathcal{M}}_{1,1}$
@@
obtained by forgetting the marked points and stabilizing. The locus of the admissible smooth bielliptic curves $\mathcal{B}_{g}^{\rm adm}$ is dense in $\overline{\mathcal{B}}_{g}^{\rm adm}$ and $\Psi:\mathcal{B}_{g}^{\rm adm}\to\mathcal{B}_{g}$ is a quasi-finite morphism obtained only by forgetting the admissible structure. It follows that
$\overline{\mathcal{B}}_{g}=\Psi(\overline{\mathcal{B}}_{g}^{\rm adm})\subseteq\overline{\mathcal{M}}_{g},$
@@
the locus of stable bielliptic curves of genus $g$, is the closure of $\mathcal{B}_{g}$ in $\overline{\mathcal{M}}_{g}$; see also Remark A.7.
The description of the boundary of $\mathcal{B}_{g}$ inside $\overline{\mathcal{B}}_{g}$ is analogous to that given in Section 2.1. We highlight some details in the following remark.
Remark 3.15.
By [SvZ20, Section 3.4], and similarly to our description in Section 2.1 and (1), the boundary of $\mathcal{B}_{g}$ in $\overline{\mathcal{B}}_{g}$ is the closure of the locus consisting of stabilizations of double covers of
$(E,O_{E})\cup_{O_{E}=\infty}(\mathbb{P}^{1},\infty)\quad\text{or}\quad E_{0}=(\mathbb{P}^{1},\infty,0,1)/\{0=1\},$ (5)
ramified at some points, where $E$ is a smooth elliptic curve and $E_{0}$ is the singular elliptic curve, i.e., singular stable $1$-pointed genus-11 curve.
In particular, $\overline{\mathcal{B}}_{g}\cap\Delta_{0}=\Gamma_{ns}^{\prime}\cup\Gamma_{s}^{\prime}$, where:
  • $\cdot$
    $\Gamma_{ns}^{\prime}$ is the closure of the locus (of stabilizations) of double covers of $(E,O_{E})\cup_{O_{E}=\infty}(\mathbb{P}^{1},\infty)$ as in (5), where $(E,O_{E})$ is a smooth elliptic curve, and there is an even number of branch points of this cover on both $E-\{O_{E}\}$ and $\mathbb{P}^{1}-\{\infty\}$. Set-theoretically, note that $\Gamma_{ns}^{\prime}$ equals the union of $\overline{\mathcal{B}}_{E,g}\cap\Delta_{0}$ over all isomorphism classes $[E]$ of smooth elliptic curves $E$.
  • $\cdot$
    $\Gamma_{s}^{\prime}$ is the closure of the locus (of stabilizations) of double covers of $E_{0}$ as in (5), where $E_{0}=(\mathbb{P}^{1},\infty,0,1)/\{0=1\}$ is the singular elliptic curve, branched at 2g22g-2 points in $\mathbb{P}^{1}-\{0,1\}$} (one of which is $\infty$) and at the node $0=1$. Note that every double cover $D\to E_{0}$ is of the form $D=(H,0^{\prime},1^{\prime})/\{0^{\prime}=1^{\prime}\}$, where $H\to\mathbb{P}^{1}$ is a smooth hyperelliptic curve of genus g1g-1, and $0^{\prime},1^{\prime}\in H$ are ramification points above $0,1\in\mathbb{P}^{1}$. Consequently, $\Gamma_{s}^{\prime}$ consists of isomorphism classes of $(H,0^{\prime},1^{\prime})/\{0^{\prime}=1^{\prime}\}$, where $H$ is a stable hyperelliptic curve of genus g1g-1, and $0^{\prime},1^{\prime}$ are two ramification points of the morphism $H\to C$, where $C$ is a rational curve.
Relying on Theorem 3.12 above, we obtain the following result on the $p$-rank stratification of the full bielliptic locus $\overline{\mathcal{B}}_{g}$.
Corollary 3.16.
For every $g\geq 2$ and $0\leq f\leq g$, the $p$-rank $\leq f$ locus of stable bielliptic curves $V_{f}(\overline{\mathcal{B}}_{g})$ is pure of dimension g2+fg-2+f.
@@
Let $W$ be a component of $V_{f}(\overline{\mathcal{B}}_{g})$. As before, we find that $\dim W\geq g-2+f$, using the purity of the $p$-rank stratification of $\overline{\mathcal{M}}_{g}$. Consider the morphism $\Phi:\overline{\mathcal{B}}_{g}^{\rm adm}\to{\overline{\mathcal{M}}}_{1,2g-2}\to{\overline{\mathcal{M}}}_{1,1}$ induced by $(D\to(E^{\prime},P_{1},\ldots,P_{2g-2}))\mapsto({E},P_{1})$, where $E=({E},P_{1})$ denotes the elliptic curve obtained by stabilizing $(E^{\prime},P_{1})$. Let $W^{\prime}$ denote the preimage of $W$ under the map $\Psi:\overline{\mathcal{B}}_{g}^{\rm adm}\to\overline{\mathcal{M}}_{g}$. We separate the proof into different cases depending on whether the generic point of $W$ is represented by a smooth curve.
  1. 1.
    Suppose that the generic point of $W$ is represented by a smooth curve; note that $\dim W=\dim W^{\prime}$. There are two possibilities:
    1. (a)
      $\Phi(W^{\prime})$ is a single point in $\mathcal{M}_{1,1}\subseteq\overline{\mathcal{M}}_{1,1}$ corresponding to a smooth elliptic curve $E$. In this case, $W$ is an irreducible locus consisting of isomorphism classes of smooth curves $D$ with $f_{D}\leq f$, each admitting a double cover $D\to E$. In other words, $W\subseteq V_{f}(\overline{\mathcal{B}}_{E,g})$. Therefore, dimW ≤dimV_f(B_E, g) = g - 2 + f - f_E ≤g - 2 + f by Theorem 3.12.
    2. (b)
      Otherwise, $\Phi(W^{\prime})$ consists of at least two points in $\overline{\mathcal{M}}_{1,1}$. Moreover, $\Phi(W^{\prime})\subseteq\overline{\mathcal{M}}_{1,1}$ is irreducible, being the image of an irreducible locus under a morphism. These two properties imply that $\Phi(W^{\prime})=\overline{\mathcal{M}}_{1,1}$. In particular, there exists an element of $W^{\prime}$ that maps to the singular elliptic curve $E_{0}=(\mathbb{P}^{1},\infty,0,1)/\{0=1\}$ under $\Phi$, and hence whose image under $\Psi$ lies in $W\cap\Delta_{0}$. Therefore, $W\cap\Delta_{0}\neq\varnothing$. Since $\overline{\mathcal{M}}_{g}$ is a smooth stack, we have $\dim W\leq\dim(W\cap\Delta_{0})+1$. Thus, to conclude that $\dim W\leq g-2+f$, it suffices to show that $\dim(W\cap\Delta_{0})\leq g-3+f$.
      Let $\Gamma_{ns}=\Gamma_{ns}^{\prime}\cap W$ and $\Gamma_{s}=\Gamma_{s}^{\prime}\cap W$, where $\Gamma_{ns}^{\prime}$ and $\Gamma_{s}^{\prime}$ are the loci described in Remark 3.15. It suffices to show that $\dim\Gamma_{ns}\leq g-3+f$ and $\dim\Gamma_{s}\leq g-3+f$. Recall that $W\subseteq V_{f}(\overline{\mathcal{B}}_{g})$.
@@
        To show $\dim\Gamma_{ns}\leq g-3+f$, note that $\Gamma_{ns}$ is the union of the loci $V_{f}(\overline{\mathcal{B}}_{E,g})\cap\Delta_{0}$ over all isomorphism classes $[E]$ of (smooth) elliptic curves $E$. By Theorem 3.12 and Lemmas 3.10 and 3.11, no component of $V_{f}(\overline{\mathcal{B}}_{E,g})$ is contained in $\Delta_{0}$, and $\dim(V_{f}(\overline{\mathcal{B}}_{E,g})\cap\Delta_{0})\leq g-3+f-f_{E}$. Moreover, the locus of curves in $\mathcal{M}_{1,1}$ with $p$-rank $f_{E}$ has dimension $f_{E}$ for $0\leq f_{E}\leq 1$. Hence, we conclude that
        $\dim\Gamma_{ns}\leq(g-3+f-f_{E})+f_{E}=g-3+f.$
      • $\cdot$
        To show $\dim\Gamma_{s}\leq g-3+f$, let $V_{f-1}(\overline{\mathcal{H}}_{g-1,2})$ be the locus of stable $2$-pointed hyperelliptic curves $(H,q_{1},q_{2})$ of genus g1g-1 whose $p$-rank is $\leq f-1$, and let $\Gamma^{\prime\prime}\subseteq V_{f-1}(\overline{\mathcal{H}}_{g-1,2})$ be the sublocus where the marked points $q_{1},q_{2}$ of $(H,q_{1},q_{2})$ are chosen to be two ramification points of the double cover $H\to C$, with $C$ a rational curve. Since there are only finitely many ramification points, we have $\dim\Gamma^{\prime\prime}=\dim V_{f-1}(\overline{\mathcal{H}}_{g-1})$. Using the definition of $\Gamma_{s}^{\prime}$ and the inclusion $W\subseteq V_{f}(\overline{\mathcal{B}}_{g})$, we deduce that $\Gamma_{s}=\Gamma_{s}^{\prime}\cap W$ is contained in the image of $\Gamma^{\prime\prime}$ under the clutching morphism $\overline{\mathcal{M}}_{g-1,2}\to\overline{\mathcal{M}}_{g}$. Lemma 3.5 then implies that
        $\dim\Gamma_{s}=\dim V_{f}(\overline{\mathcal{H}}_{g-1})\leq g-3+f.$
@@
    If $W$ is not generically represented by a smooth curve, then $W$ consists entirely of isomorphism classes of singular curves. If the image of $W^{\prime}$ under $\Phi$ contains the isomorphism class of $E_{0}=(\mathbb{P}^{1},\infty,0,1)/\{0=1\}$, the same discussion as above implies that $W\cap\Delta_{0}\neq\varnothing$ and $\dim W\leq g-2+f$; this covers the cases $\Phi(W^{\prime})=\overline{\mathcal{M}}_{1,1}$ or $\Phi(W^{\prime})=[E_{0}]$. Otherwise, $\Phi(W^{\prime})=[E]$, where $E=(E,O_{E})$ is a smooth elliptic curve. As before, Theorem 3.12 implies that $\dim W\leq\dim V_{f}(\overline{\mathcal{B}}_{E,g})=g-2+f-f_{E}\leq g-2+f$. (More precisely, $\dim W\geq g-2+f$, and Lemmas 3.10 and 3.11 imply that $f_{E}=0$ in this case and that $W$ could only possibly be the closure of the locus consisting of isomorphism classes of double covers of curves of the form $(E,O_{E})\cup_{O_{E}=\infty}(\mathbb{P}^{1},\infty)$, ramified above 2g22g-2 distinct points $P_{i}\in\mathbb{P}^{1}-\{\infty\}$, $1\leq i\leq 2g-2$, and unramified otherwise, as in Corollary 3.14.)